OpenStudy (jackfrost33):

HOW WOULD YOU WRITE THE SLOPE INTERCEPT EQUATION OF THE LINE THAT GOES THROUGH THE POINTS (-5, 2) AND (-3,-2)

OpenStudy (jackfrost33):

@zepdrix help please???

OpenStudy (ilovemath14):

The links might help you and if they don't tag me

OpenStudy (mtalhahassan2):

m= y2-y1/ x2-x1 =-2-2/ -3+5 =-4/2 m =-2 would be the slope

OpenStudy (mtalhahassan2):

y=mx+b are you familiar with this equation?

OpenStudy (jackfrost33):

yes i am but my brother is not and i can not explain it to him correctly so therfore i am of no use

OpenStudy (mtalhahassan2):

so we already find the m now we need to find b which is y-intercept

OpenStudy (jackfrost33):

how would you do that? ~her brother~

OpenStudy (mtalhahassan2):

ok first go and watch that video you will understand better

OpenStudy (ilovemath14):

Look at the links i put they will help you

OpenStudy (jackfrost33):

my teacher tried to help me by showing me that video but it was over my head

OpenStudy (jackfrost33):

but ill just ask her on monday thanks anyway

OpenStudy (3mar):

May I help?

OpenStudy (jackfrost33):

yes you may

OpenStudy (3mar):

Ok. Thanks.

OpenStudy (jackfrost33):

anything is appreciated

OpenStudy (3mar):

The general form of line equation passes through two points is:

OpenStudy (jackfrost33):

my teacher said that it would be a positive answer but i got negitive 2

OpenStudy (3mar):

Let's do it together?

OpenStudy (3mar):

Maths don't tell lie

OpenStudy (jackfrost33):

im pretty sure i did it wrong

OpenStudy (3mar):

No problem. What matters is you get the right answer!

OpenStudy (3mar):

Let's

OpenStudy (jackfrost33):

yeh it helps to get it right lols

OpenStudy (3mar):

\[y_2-y_1=?\]

OpenStudy (3mar):

hard?

OpenStudy (3mar):

Do you follow?

OpenStudy (3mar):

Hey